# LoSA: Long-Short-range Adapter for Scaling End-to-End Temporal Action Localization

### v1

@@ Abstract
Temporal Action Localization (TAL) involves localizing and classifying action snippets in an untrimmed video. The emergence of large video foundation models has led RGB-only video backbones to outperform previous methods needing both RGB and optical flow modalities. Leveraging these large models is often limited to training only the TAL head due to the prohibitively large GPU memory required to adapt the video backbone for TAL. To overcome this limitation, we introduce LoSA, the first memory-and-parameter-efficient backbone adapter designed specifically for TAL to handle untrimmed videos. LoSA specializes for TAL by introducing Long-Short-range Adapters that adapt the intermediate layers of the video backbone over different temporal ranges. These adapters run parallel to the video backbone to significantly reduce memory footprint. LoSA also includes Long-Short-range Fusion that strategically combines the output of these adapters from the video backbone layers to enhance the video features provided to the TAL head. Experiments show that LoSA significantly outperforms all existing methods on standard TAL benchmarks, THUMOS-14 and ActivityNet-v1.3, by scaling end-to-end backbone adaptation to billion-parameter-plus models like VideoMAEv2 (ViT-g) and leveraging them beyond head-only transfer learning.
* Equal Contribution
This work was done as Akshita’s internship project at Microsoft.
@@ 1 Introduction
Temporal Action Localization (TAL) refers to localizing and classifying action snippets in an untrimmed (arbitrarily long) video. TAL is crucial for applications in video indexing/search, surveillance, responsible AI, and robotics [37]. Several methods have emerged for TAL over the years that formulate it as a downstream transfer learning task [44, 28, 34, 39]. Most works [44, 20] perform head-only transfer learning where a frozen video backbone, generally pretrained on a large corpus of trimmed ($<$30s) videos like Kinetics-600 [6], is employed to extract features from untrimmed videos (Fig. 1a). These features are then concatenated and fed to a trainable head designed to perform TAL. In this context, while certain studies [44, 29, 31] have shown improved results using both RGB and optical flow features, advances in video foundation models have enabled recent works [34, 36] to employ models like ViT-g with over 1 B parameters [43] to surpass previous methods with RGB features alone. This is because the effectiveness of data and model scaling is able to offset the need for expensive optical flow estimation (Fig 1e).
@@
More recently, Parameter-efficient Transfer Learning/Fine-Tuning (PETL/ PEFT) approaches [16, 17, 30, 42] emerged, motivated by the computational constraints of adapting billion-parameter foundation models for downstream tasks. However, existing approaches do not have a mechanism to adapt the intermediate layers of the base foundation model to handle untrimmed videos (as showing in Fig 1c). These approaches are ill-equipped to work optimally for TAL, since understanding long-range and short-range temporal dependencies is required to correctly localize a wide range of actions.
To overcome this challenge, we introduce LoSA, the first memory- and-parameter-efficient backbone adapter that is tailored for TAL and untrimmed videos to harness large video foundation models more effectively beyond head-only transfer learning (Fig 1d). LoSA comprises a series of lightweight Long-range and Short-range Adapters that are attached to the intermediate layers of the video backbone. With video being processed by the large video foundation model as a sequence of trimmed video clips, these adapters learn to adapt the intermediate layers of the video backbone by capturing long-term and short-term dependencies among the video frames respectively. This allows an improved long-range temporal learning of the untrimmed video at each intermediate layer (Fig 1d) while also capturing the fine-grained short-term temporal changes in the video, allowing for more effective localization of actions. To allow each intermediate layer to contribute directly towards improving TAL, the adapters leverage cross-attention between the intermediate layers and the last layer of the video backbone to transform the output of intermediate layers.
The Long-range and Short-range Adapters run parallel to the video backbone and their outputs directly aggregate with the last layer features. This circumvents gradient backpropagation through the video backbone, which significantly reduces the memory footprint of adapting the backbone for TAL. To facilitate this aggregation, LoSA introduces Long-Short-range Fusion with a learnable gating function to weigh the contribution of each intermediate layer and fuse them together with the output of the last layer to generate improved features for the TAL head. These TAL-enhanced features enable more accurate action boundaries compared to head-only transfer learning as evident from the superior performance in Fig 1e.
@@
    LoSA comprises a novel adapter design to enable memory-and-parameter efficiency for untrimmed videos by employing a series of lightweight Long-range and Short-range adapters that run parallel to the video backbone and a Long-Short-range Fusion module to adaptively fuse the outputs from the adapters to improve TAL.
@@
    LoSA is capable of end-to-end backbone adaptation of over-billion-parameter video models beyond head-only transfer learning, establishing new SOTA for TAL.
@@ 2 Related Work
2.1 Temporal Action Localization (TAL)
Most TAL approaches leverage RGB and optical flow features pre-extracted from a video backbone. Among these are two-stage methods [39, 3, 22, 21, 14, 46], which generate pre-defined action proposals and then classify them into action classes while regressing the actual action boundaries, and one-stage methods [20, 24, 44, 29, 31], which perform TAL in a single pass without separately generating action proposals. These approaches perform head-only transfer learning, treating the video backbone as frozen. In spite of shallow training and leveraging relatively small backbones like I3D [7], TSN [35], and TSP [1], they achieve competitive performance by using optical flow that enhances temporal understanding. However, optical flow estimation is computationally expensive, making it challenging to scale on increasingly large video datasets. Recently, large video foundation models [34, 36] have demonstrated superior performance on TAL using RGB features only. These are also limited to head-only transfer learning due to the prohibitively large GPU memory footprint for end-to-end training.
2.2 Backbone Adaptation approaches for TAL
There exist approaches [8, 23, 45, 9] that attempt to adapt an RGB-only video backbone beyond head-only transfer learning to mitigate the need for optical flow. They do so via memory optimizations such as reducing spatial resolution [23], channel activations [9], feature caching [8], and rewiring the backbone [45]. While they can operate on relatively small backbones like SlowFast-101 [13], ViT-B [12], and ResNet-50 [15], they fail to scale to the size of current visual foundation models with potentially billions of parameters [34]. LoSA, with its memory- and parameter-efficient backbone adapter, mitigates this issue and enables backbone adaptation to leverage RGB-only large video backbones beyond head-only transfer learning to outperform all existing methods.
2.3 Parameter-efficient Transfer Learning (PETL)
With the advent of large-language models (LLMs) [4, 32], parameter-efficient transfer learning/finetuning (PETL/PEFT) [17, 16] has emerged to reduce computational costs of finetuning LLMs on downstream tasks. Inspired by LLM-based PETL, vision-based PETL was developed to enable efficient transfer learning on visual tasks. Yet, most approaches are parameter-efficient but not memory-efficient [30, 40] as their design causes gradient backpropagation through the backbone. Some more recent methods [41, 30, 11] attempt to address memory efficiency, but no existing method, to the best of our knowledge, is suited to handle untrimmed videos. LoSA is the first memory- and parameter-efficient approach that is designed for TAL.
@@ 3 Method
Figure 2: LoSA Overview: LoSA comprises a series of Long-range and Short-range Adapters that attach to the intermediate layers $11,\ldots,N-$ 1 of a video backbone. (a) Each Short-range Adapter consists of a cross-attention module that uses the video clip-level spatio-temporal features of an intermediate layer as Query and the last layer temporally concatenated features as Key and Value. (b) Similarly, each Long-range Adapter uses a cross-attention module to cross-attend the temporally concatenated long-range untrimmed video features of an intermediate layer as Query (Q) and the last layer temporally concatenated features as Key (K) and Value (V). (c) Finally, the Long-Short-range Fusion module learns scaling parameters to gate the contribution of the Long-range and Short-range Adapters and combines them with the temporally concatenated last layer features via a projection layer to generate the TAL-enhanced features going into the learnable TAL head for outputting the localized action snippets.
We describe the different components of LoSA in the following subsections, starting with laying out the preliminaries. Fig 2 provides an overview of the model.
@@ 3.1 Preliminaries
Let $g$ denote a video backbone comprising $N$ layers, $f_{1},\ldots,f_{N}$, defined as $g=f_{N}(f_{N-1}(\ldots f_{1}(\mathbf{X})\ldots))$, where $\mathbf{X}$ is the input to the backbone. Let $F_{i}$ be the feature representation obtained as the output from layer $f_{i}$.
To adapt $g$ for TAL, $\mathbf{X}$ is an untrimmed video comprising an arbitrary number of frames. Since the existing video backbones are trained using trimmed video clips, we divide $\mathbf{X}$ into a sequence of $T$ clips ${x_{1},\ldots,x_{T}}$ where each clip is a sequence of $T^{\prime}$ frames such that $T^{\prime}\ll$ total number of frames in the untrimmed video. Depending on the stride, the clips can be either overlapping or disjoint. We feed each clip $x_{t}~{}\forall~{}t\in\{1\ldots T\}$ to the video backbone $g$ to generate a set of feature maps $F_{i}^{x_{t}}~{}\forall~{}i\in\{1\ldots N\},t\in\{1\ldots T\}$ corresponding to all the layers of $g$. We assume each feature map $F_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ where $T_{i}$, $H_{i}$, $W_{i}$, and $C_{i}$ denote the temporal, height, width, and channel dimensions of the video clip features at intemediate layer $f_{i}$. Further, at each intermediate layer $f_{i}$, let $\mathbf{F^{X}_{i}}=\{F_{i}^{x_{1}},\ldots,F_{i}^{x_{T}}\}\in\mathbb{R}^{TT_{i}\times{H_{i}\times W_{i}\times C_{i}}}$ be the concatenation of the feature maps along the temporal dimension.
@@ 3.2 Short-range Temporal Adapter
Let us consider an intermediate spatio-temporal feature map $F_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ from layer $f_{i}$ at temporal location $t$ where $i\in\{1\ldots N-1\}$ and $t\in\{1\ldots T\}$ obtained from feeding a trimmed video clip of the untrimmed video to the video backbone. Since $T^{\prime}\ll$ total number of frames in the untrimmed video, $F_{i}^{x_{t}}$ captures a short-range temporal context and provides a fine-grained temporal understanding in the local temporal neighborhood of the untrimmed video around its timestamp $t$. The video backbone processes each $F_{i}^{x_{t}}~{}\forall~{}t\in\{1\ldots T\}$ independent of each other, which prevents the context of the full untrimmed video to influence the representation learning. So to perform transfer learning over these features for effective end-to-end action localization, LoSA comprises a series of Short-range Temporal Adapters (Fig 2a) to adapt $F_{i}^{x_{t}}~{}\forall~{}t\in\{1\ldots T\},i\in\{1\ldots N-1\}$ conditioned on all the short-range spatio-temporal feature maps from the last layer, $F_{N}^{x_{t}}~{}\forall~{}t\in\{1\ldots T\}$.
For this, we introduce a multi-headed cross-attention module in the Short-range Temporal Adapter at each intermediate layer $f_{i}~{}\forall~{}i\in\{1\ldots N-1\}$ of the video backbone. We first employ a convolutional block at each layer $i\in\{1\ldots N\}$ of the video backbone to process the spatial dimensions by transforming $F_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}\rightarrow{F^{\prime}}_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times C_{i}}$. We then feed ${F^{\prime}}_{i}^{x_{t}}$ as Query and $\mathbf{{F^{\prime}}^{X}_{N}}$ as Key and Value to the cross-attention module, where $\mathbf{{F^{\prime}}^{X}_{N}}=\{{F^{\prime}}_{N}^{x_{1}},\ldots,{F^{\prime}}_{N}^{x_{T}}\}$ is the concatenation of the transformed feature maps from last layer $N$ of the video backbone along the temporal dimension. The cross-attention module outputs $FS^{x_{t}}_{i}\in\mathbb{R}^{T_{i}\times C_{i}}~{}\forall~{}t\in\{1\ldots T\},i\in\{1\ldots N-1\}$, representing the short-range features adapted for end-to-end TAL.
@@ 3.3 Long-range Temporal Adapter
Similar to the Short-range Temporal Adapters, the Long-range Temporal Adapter adapts the feature map sequence at a given intermediate layer conditioned on the feature map sequence from the last layer of the video backbone. For this, each Long-range Temporal Adapter consists of a cross-attention module at each intermediate layer $f_{i}~{}\forall~{}i\in\{1\ldots N-1\}$ of the video backbone. We feed the output of the convolutional block, concatenated along temporal dimension, $\mathbf{{F^{\prime}}^{X}_{i}}$ as Query and $\mathbf{{F^{\prime}}^{X}_{N}}$ as Key and Value to the cross-attention module. By doing so, we consider the last layer features as reference to cross-attend the intermediate long-range temporal features to adapt the latter into what directly improves the features going into TAL. The cross-attention module outputs $\mathbf{FL^{X}_{i}}\in\mathbb{R}^{TT_{i}\times C_{i}}~{}\forall~{}i\in\{1\ldots N-1\}$, representing the long-range features adapted for end-to-end TAL.
3.4 Long-Short-range Fusion

### v2

@@ Abstract
Temporal Action Localization (TAL) involves localizing and classifying action snippets in an untrimmed video. The emergence of large video foundation models has led RGB-only video backbones to outperform previous methods needing both RGB and optical flow modalities. Leveraging these large models is often limited to training only the TAL head due to the prohibitively large GPU memory required to adapt the video backbone for TAL. To overcome this limitation, we introduce LoSA, the first memory-and-parameter-efficient backbone adapter designed specifically for TAL to handle untrimmed videos. LoSA specializes for TAL by introducing Long-Short-range Adapters that adapt the intermediate layers of the video backbone over different temporal ranges. These adapters run parallel to the video backbone to significantly reduce memory footprint. LoSA also includes Long-Short-range Gated Fusion that strategically combines the output of these adapters from the video backbone layers to enhance the video features provided to the TAL head. Experiments show that LoSA significantly outperforms all existing methods on standard TAL benchmarks, THUMOS-14 and ActivityNet-v1.3, by scaling end-to-end backbone adaptation to billion-parameter-plus models like VideoMAEv2 (ViT-g) and leveraging them beyond head-only transfer learning.
* Equal Contribution. This work was done as Akshita’s internship project at Microsoft.
@@ 1 Introduction
Temporal Action Localization (TAL) refers to localizing and classifying action snippets in an untrimmed (arbitrarily long) video. TAL is crucial for applications in video indexing/search, surveillance, responsible AI, and robotics [36]. Many TAL methods treat it as a downstream transfer learning task [43, 27, 33, 38]. Most works [43, 19] perform head-only transfer learning where a frozen video backbone, generally pretrained on a large corpus of trimmed ($<$30s) videos like Kinetics-600 [6], is employed to extract features from untrimmed videos (Fig 1a). These features are then concatenated and fed to a trainable head designed to perform TAL. In this context, while certain studies [43, 28, 30] have shown improved results using both RGB and optical flow features, advances in video foundation models have enabled recent works [33, 35] to employ models like ViT-g with over 1 B parameters [42] to surpass previous methods with RGB features alone. This is because the effectiveness of data and model scaling is able to offset the need for expensive optical flow estimation (Fig 1e).
@@
Recently, Parameter-efficient Transfer Learning/Fine-Tuning (PETL/ PEFT) approaches [15, 16, 29, 41] emerged, motivated by the computational constraints of adapting billion-parameter foundation models for downstream tasks. However, existing approaches are ill-equipped to learn context in untrimmed videos over different temporal ranges (Fig 1c) which is crucial to correctly localize actions of diverse type and duration [17, 5]. These methods are thus sub-optimal in adapting the base foundation model for TAL.
To overcome this challenge, we introduce LoSA, the first memory- and-parameter-efficient backbone adapter that is tailored for TAL and untrimmed videos to harness large video foundation models more effectively beyond head-only transfer learning (Fig 1d). LoSA comprises a series of lightweight Long-range and Short-range Adapters that are attached to the intermediate layers of the video backbone. With video being processed by the large video foundation model as a sequence of trimmed video clips, these adapters learn to adapt the intermediate layers of the video backbone by capturing long-term and short-term dependencies among the video frames respectively. This allows an improved long-range temporal learning of the untrimmed video at each intermediate layer (Fig 1d) while also capturing the fine-grained short-term temporal changes in the video, allowing for more effective localization of actions. To allow each intermediate layer to contribute directly towards improving TAL, the adapters leverage cross-attention between intermediate layers and the last layer of the video backbone to transform the output of intermediate layers.
The Long-range and Short-range Adapters run parallel to the video backbone and their outputs directly aggregate with the last layer features. This circumvents gradient backpropagation through the video backbone, which significantly reduces the memory footprint of adapting the backbone for TAL. To facilitate this aggregation, LoSA introduces Long-Short-range Gated Fusion with a learnable gating function to weigh the contribution of each intermediate layer and fuse them together with the output of the last layer to generate improved features for the TAL head. These TAL-enhanced features enable more accurate action boundaries compared to head-only transfer learning as evident from the superior performance in Fig 1e.
@@
    LoSA comprises a novel adapter design to enable memory-and-parameter efficiency for untrimmed videos by employing a series of lightweight Long-range and Short-range adapters that run parallel to the video backbone and a Long-Short-range Gated Fusion module to adaptively fuse the outputs from the adapters to improve TAL.
@@
    LoSA is capable of end-to-end backbone adaptation of $>$1 B parameter video models beyond head-only transfer learning, establishing new SOTA for TAL.
@@ 2 Related Work
Temporal Action Localization (TAL). Most TAL approaches leverage RGB and optical flow features pre-extracted from a video backbone. Among these are two-stage methods [38, 3, 21, 20, 13, 45], which generate pre-defined action proposals and then classify them into action classes while regressing the actual action boundaries, and one-stage methods [19, 23, 43, 28, 30], which perform TAL in a single pass without separately generating action proposals. These approaches perform head-only transfer learning, treating the video backbone as frozen. In spite of shallow training and leveraging relatively small backbones like I3D [7], TSN [34], and TSP [1], they achieve competitive performance by using optical flow that enhances temporal understanding. However, optical flow estimation is computationally expensive, making it challenging to scale on increasingly large video datasets. Recently, large video foundation models [33, 35] have demonstrated superior performance on TAL using RGB features only. These are also limited to head-only transfer learning due to the prohibitively large GPU memory footprint for end-to-end training.
Backbone Adaptation approaches for TAL. There exist approaches [8, 22, 44, 9] that attempt to adapt an RGB-only video backbone beyond head-only transfer learning to mitigate the need for optical flow. They do so via memory optimizations such as reducing spatial resolution [22], channel activations [9], feature caching [8], and rewiring the backbone [44]. While they can operate on relatively small backbones like SlowFast-101 [12], ViT-B [11], and ResNet-50 [14], they fail to scale to the size of current visual foundation models with billions of parameters [33]. LoSA, with its memory- and parameter-efficient backbone adapter, mitigates this issue and enables backbone adaptation of RGB-only large video backbones beyond head-only transfer learning to outperform all existing methods.
Parameter-efficient Transfer Learning (PETL). With the advent of large-language models (LLMs) [4, 31], parameter-efficient transfer learning/finetuning (PETL/PEFT) [16, 15] has emerged to reduce computational costs of finetuning LLMs on downstream tasks. Inspired by LLM-based PETL, vision-based PETL was developed to enable efficient transfer learning on visual tasks. Yet, most approaches are parameter-efficient but not memory-efficient [29, 39] as their design causes gradient backpropagation through the backbone. Some more recent methods [40, 29] attempt to address memory efficiency, but no existing method, to the best of our knowledge, is suited to handle untrimmed videos. LoSA is the first memory- and parameter-efficient approach that is designed for TAL.
@@ 3 Method
Figure 2: LoSA Overview: LoSA comprises a series of Long-range and Short-range Adapters that attach to the intermediate layers $11,\ldots,N-$ 1 of a video backbone. (a) Each Short-range Adapter consists of a cross-attention module that uses the video clip-level spatio-temporal features of an intermediate layer as Query and the last layer temporally concatenated features as Key and Value. (b) Similarly, each Long-range Adapter uses a cross-attention module to cross-attend the temporally concatenated long-range untrimmed video features of an intermediate layer as Query (Q) and the last layer temporally concatenated features as Key (K) and Value (V). (c) Finally, the Long-Short-range Gated Fusion module learns scaling parameters to gate the contribution of the Long-range and Short-range Adapters and combines them with the temporally concatenated last layer features via a projection layer to generate the TAL-enhanced features going into the learnable TAL head for outputting the localized action snippets.
We describe the components of LoSA in the following subsections. Fig 2 provides an overview of the model.
@@ 3.1 Preliminaries
Let $g$ denote a video backbone comprising $N$ layers, $f_{1},\ldots,f_{N}$, defined as $g=f_{N}(f_{N-1}(\ldots f_{1}(\mathbf{X})\ldots))$, where $\mathbf{X}$ is the input to the backbone. Let $F_{i}$ be the feature representation obtained as the output from layer $f_{i}$. To adapt $g$ for TAL, $\mathbf{X}$ is an untrimmed video comprising an arbitrary number of frames. Since the existing video backbones are trained using trimmed video clips, we divide $\mathbf{X}$ into a sequence of $T$ clips ${x_{1},\ldots,x_{T}}$ where each clip is a sequence of $T^{\prime}$ frames such that $T^{\prime}\ll$ total number of frames in the untrimmed video. Depending on the stride, the clips can be either overlapping or disjoint. We feed each clip $x_{t}~{}\forall~{}t\in\{1\ldots T\}$ to the video backbone $g$ to generate a set of feature maps $F_{i}^{x_{t}}~{}\forall~{}i\in\{1\ldots N\},t\in\{1\ldots T\}$ corresponding to all the layers of $g$. We assume each feature map $F_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ where $T_{i}$, $H_{i}$, $W_{i}$, and $C_{i}$ denote the temporal, height, width, and channel dimensions of the video clip features at intemediate layer $f_{i}$. Further, at each intermediate layer $f_{i}$, let $\mathbf{F^{X}_{i}}=\{F_{i}^{x_{1}},\ldots,F_{i}^{x_{T}}\}\in\mathbb{R}^{TT_{i}\times{H_{i}\times W_{i}\times C_{i}}}$ be the concatenation of the feature maps along the temporal dimension.
@@ 3.2 Short-range Temporal Adapter
Let $F_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}$ be an intermediate spatio-temporal feature map from layer $f_{i}$ at temporal location $t$ where $i\in\{1\ldots N-1\}$ and $t\in\{1\ldots T\}$ obtained from feeding a trimmed video clip of the untrimmed video to the video backbone. Since $T^{\prime}\ll$ total number of frames in the untrimmed video, $F_{i}^{x_{t}}$ captures a short-range temporal context and provides a fine-grained temporal understanding in the local temporal neighborhood of the untrimmed video around its timestamp $t$. The video backbone processes each $F_{i}^{x_{t}}~{}\forall~{}t\in\{1\ldots T\}$ independent of each other, which prevents the context of the full untrimmed video to influence the representation learning. So to perform transfer learning over these features for effective end-to-end TAL, LoSA comprises a series of Short-range Temporal Adapters (Fig 2a) to adapt $F_{i}^{x_{t}}~{}\forall~{}t\in\{1\ldots T\},i\in\{1\ldots N-1\}$ conditioned on all the short-range spatio-temporal features from the last layer, $F_{N}^{x_{t}}~{}\forall~{}t\in\{1\ldots T\}$.
For this, we introduce a multi-headed cross-attention module in the Short-range Temporal Adapter at each intermediate layer $f_{i}~{}\forall~{}i\in\{1\ldots N-1\}$ of the video backbone. We first employ a convolutional block at each layer $i\in\{1\ldots N\}$ of the video backbone to process the spatial dimensions by transforming $F_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times H_{i}\times W_{i}\times C_{i}}\rightarrow{F^{\prime}}_{i}^{x_{t}}\in\mathbb{R}^{T_{i}\times C_{i}}$. For the cross-attention, we use ${F^{\prime}}_{i}^{x_{t}}$ as Query and the feature set $\{{F^{\prime}}_{N}^{x_{1}},\ldots,{F^{\prime}}_{N}^{x_{T}}\}$ from last layer $N$ as Key and Value. The cross-attention module outputs $FS^{x_{t}}_{i}\in\mathbb{R}^{T_{i}\times C_{i}}~{}\forall~{}t\in\{1\ldots T\},i\in\{1\ldots N-1\}$, denoting the short-range features adapted for end-to-end TAL.
@@ 3.3 Long-range Temporal Adapter
The Long-range Temporal Adapter adapts the feature sequence at a given intermediate layer conditioned on the feature sequence from last layer of the video backbone. For this, each Long-range Temporal Adapter consists of a cross-attention at each intermediate layer $f_{i}~{}\forall~{}i\in\{1\ldots N-1\}$ of the video backbone. We feed the output of the convolutional block, concatenated along temporal dimension, $\mathbf{{F^{\prime}}^{X}_{i}}$ as Query and $\mathbf{{F^{\prime}}^{X}_{N}}$ as Key and Value to the cross-attention module. By doing so, we consider the last layer features as reference to cross-attend the intermediate long-range temporal features to adapt the latter into what directly improves the features used for TAL. The cross-attention module outputs $\mathbf{FL^{X}_{i}}\in\mathbb{R}^{TT_{i}\times C_{i}}~{}\forall~{}i\in\{1\ldots N-1\}$, representing the long-range features adapted for end-to-end TAL.
3.4 Long-Short-range Gated Fusion